A Bibliography of Supercomputing ’96

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.07

Abstract

This bibliography records articles presented at the Supercomputing ’96 conference.

Title word cross-reference

k [PV96]. n [PV96, HJ96]. O(N) [HJ96].

-ary [PV96]. -body [HJ96]. -cubes [PV96]. -way [KK96].

4 [HLT96].

’96 [ACM96].

Adaption [BOS96]. Adaptive [PV96]. Algebra [BCC^{+}96]. Algorithm [HLJJ^{+}96]. Algorithms [KCM96]. All-optical [YMG96]. Analysis [GAL96].

Analytical [SG96]. Application [BWF^{+}96]. Application-Level [BWF^{+}96]. Applications [VL96]. Architecture [HLT96]. Array [CWwK^{+}96]. ary [PV96]. Assignment [WSC96]. Assimilation [DF96b]. Association [ZOPL96]. Atmospheric [LF96].

Balancing [BOS96]. Bandwidth [dL96]. Based [DF96a]. Benchmark [HLT96]. Beyond [Hat96]. body [HJ96]. Boundary [GKS96]. Bringing [FKKC96]. Building [DF96a].

C3I [JMV^{+}96]. Cascade [NIY^{+}96]. Case [GMW96]. Cell [ADHF96]. Circulation [LF96]. Climate [DF96b]. Clusters [CSC96]. Code [BXC96]. Codes [ADHF96].

Communication [EK96]. Compiled [YMG96]. Compiler [VL96]. Compiler-directed [VL96]. Complex [dL96]. Computation [HKW96]. Computational [BL96]. Computers [BBG^{+}96]. Computing [CHPW96]. Conference [ACM96]. Contact [PHA^{+}96]. Control [GAL96]. Cosmological [BXC96]. Counters [ZLTI96]. Cray [Hat96]. cubes [PV96]. Curved [KCM96].

Data [DF96b]. Data-Parallel [HJ96]. Design [CHPW96]. Designing [Meh96]. Detection [PHA^{+}96]. Dimensional [NIY^{+}96]. directed [VL96]. Discrete [CHPW96]. Disk [PC96]. Distributed [BWF^{+}96]. Distributed-Memory [BOS96]. Distribution [GAL96]. Double [FKKC96]. DSM [HKW96]. Dynamic [HKW96]. Dynamics [BL96].

Earthquake [BBG^{+}96]. Education [Meh96]. Effective [dL96]. Effectively [Meh96]. Element [EK96]. Element-Fourier [EK96]. Elliptic [SS96]. Empirical [WSC96]. Enabling [PC96]. Estimation [CSA96]. Evaluation [WRH96].

Factorization [FY96]. Failures [PC96]. File [PSS96]. Fine [HSS96]. Fine-Grained [HSS96]. Flow [GAL96]. Fluid [Hat96]. Fortran [ADHF96]. Fourier [EK96].

General [LF96]. Gigaflops [Hat96]. Global [BOS96]. Grained [HSS96]. Graphs [KK96]. Ground [BBG^{+}96].

Heterogeneous [BWF^{+}96]. Hierarchical [HJ96]. High [CHPW96, ADHF96].

High-performance [FGKT96]. Highly [TWL96]. HINT [SG96]. HPCC [DF96a]. HPF [FKKC96]. Hydrodynamics [BXC96].

I/O [CWwK^{+}96]. IBM [CCHv96]. Impact [SGS96]. Imperfectly [KP96]. Implementation [HJ96]. Increasing [dL96]. Injection [PV96]. Instrumentation [WRH96]. intensive [PSS96]. Interactive [KCM96]. Interface [FKKC96]. Interfaces [PSS96]. Introduction [JMV^{+}96]. Irregular [KK96]. Issues [BCC^{+}96]. Iterative [VL96].

Job [SGS96].

Language [HSS96]. Large [CHPW96]. Latency [CCHv96]. Level [BWF^{+}96]. Library [CWwK^{+}96]. Lightweight [BL96]. Limited [PV96]. Linear [BCC^{+}96]. Load [BOS96]. Loops [KP96]. Low [CCHv96]. Low-Latency [CCHv96]. LU [FY96].

Machine [DF96a]. Machines [FY96]. Massively [DF96b]. Materials [Meh96]. Matrix [HLJJ^{+}96]. memory [VL96]. Memory-intensive [PSS96]. Mesh [BOS96]. Message [FKKC96]. Metacomputing [FGKT96]. Methods [HJ96]. Metric [SG96]. Migration [HKW96]. Minimal [PV96]. Mining [ZOPL96]. MIPS [ZLTI96]. Mission [TWL96]. Mix [SGS96]. Model [SG96]. Modeling [BBG^{+}96]. Models [EK96]. Molecular [BL96]. Motion [BBG^{+}96]. MPI [OM96]. Multi [ZOPL96]. Multi-processors [ZOPL96]. Multilevel [KK96]. Multimethod [FGKT96]. Multiplication [HLJJ^{+}96]. Multivector [dL96].

Navier [EK96]. NCAR [HLT96]. NEC [HLT96]. Nested [KP96]. NetSolve [CD96]. Network [CD96]. Networks [BWF^{+}96]. November [ACM96]. Numerical [NIY^{+}96]. NWT [NIY^{+}96].

O [CWwK^{+}96]. OMPI [OM96]. optical [YMG96]. Optimization [LF96]. Optimizations [SGS96]. Optimizing [OM96].

PA [ACM96]. Panda [CWwK^{+}96]. Paradyn [WRH96]. Parallel [JMV^{+}96]. Parallelism [FKKC96]. Partial [OM96]. Particle [ADHF96]. Particle-in-Cell [ADHF96]. Partitioning [KK96]. Passing [FKKC96]. Patterns [EK96]. PDEs [SS96]. Performance [SG96]. Pittsburgh [ACM96]. Pivoting [FY96]. Planetary [TWL96]. Portable [GMW96]. Preconditioners [GKS96]. Preliminary [JMV^{+}96]. Prism [EK96]. Problems [CD96]. Proceedings [ACM96]. Processors [dL96]. Profiling [HSS96]. Programs [OM96]. Properties [BCG^{+}96]. PVM [CSC96].

R10000 [ZLTI96]. Renderer [TWL96]. Results [JMV^{+}96]. Rheological [BCG^{+}96]. RISC [CCHv96]. RM2 [PC96]. Routing [PV96]. Rules [ZOPL96]. Runtime [WSC96].

Scalable [KCM96]. ScaLAPACK [BCC^{+}96]. Scale [CHPW96]. Schedulers [SGS96]. Scheduling [BWF^{+}96]. Scheme [KK96]. Science [CD96]. Scientific [GMW96]. Server [CD96]. Shared [VL96]. Shared-memory [VL96]. Sharing [SGS96]. Simulation [BCG^{+}96]. Simulations [BL96]. Smoothed [PHA^{+}96]. SnB [GMW96]. Soil [CHPW96].

Solver [EK96]. Solvers [GKS96]. Solving [CD96]. SP [CCHv96]. Space [SGS96]. Sparse [FY96]. Spectral [EK96]. Standards [FKKC96]. Steering [BL96]. Stereo [TWL96]. Stokes [EK96]. Strassen [HLJJ^{+}96]. STREN [TWL96]. Striping [PC96]. Structure [CSA96]. Study [GMW96]. Suite [HLT96]. Supercomputer [DF96b]. Supercomputers [BCG^{+}96]. Supercomputing [ACM96]. Surfaces [KCM96]. SX [HLT96]. SX-4 [HLT96]. System [CCHv96, WRH96]. System/6000 [CCHv96]. Systems [CHPW96].

Task [FKKC96]. TDM [YMG96].

Technical [Meh96]. Technologies [DF96a]. Terrain [TWL96]. Testing [WRH96]. Tolerance [PC96]. Tool [GMW96]. Toroidal [PV96]. Transformations [KP96]. Transient [PHA^{+}96]. Tunnel [NIY^{+}96].

UCLA [LF96]. Using [GMW96].

Versus [PSS96]. Very [BL96]. Via [FKKC96]. Virtual [DF96a]. Visualization [KCM96].

way [KK96]. Web [DF96a]. Wide [DF96a]. Wind [NIY^{+}96]. World [DF96a]. World-Wide [DF96a]. WWW [Meh96].

References

ACM:1996:SCP

[ACM96] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17–22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN ????

URL http://www.supercomp.
org/sc96/proceedings/. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Akarsu:1996:PCS

[ADHF96] Erol Akarsu, Kivanc Dincer, Tomasz Haupt, and Geoffrey C. Fox. Particle-in-cell simulation codes in High Performance Fortran. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
AKARSU/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Bao:1996:EGM

[BBG^{+}96] Hesheng Bao, Jacobo Bielak, Omar Ghattas, Loukas F. Kallivokas, David R. O’Hallaron, Jonathan R. Shewchuk, and Jifeng Xu. Earthquake ground motion modeling on parallel computers. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
OGHATTAS/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Blackford:1996:SPL

[BCC^{+}96] Laura Susan Blackford, J. Choi, A. Cleary, A. Petitet, R. C. Whaley, J. Demmel, I. Dhillon, K. Stanley, J. Dongarra, S. Hammarling, G. Henry, and D. Walker. ScaLAPACK: A portable linear algebra library for distributed memory computers — design issues and performance. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
DONGARRA/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Bhupathiraju:1996:MSR

[BCG^{+}96] R. K. Bhupathiraju, S. T. Cui, S. Gupta, H. D. Cochran, and P. T. Cummings. Molecular simulation of rheological properties using massively parallel supercomputers. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
CUMMINGS/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Beazley:1996:LCS

[BL96] David M. Beazley and Peter S. Lomdahl. Lightweight computational steering of very large scale molecular dynamics simulations. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
BEAZLEY/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Biswas:1996:GLB

[BOS96] Rupak Biswas, Leonid Oliker, and Andrew Sohn. Global load balancing with parallel mesh adaption on distributedmemory systems. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
BISWAS/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Berman:1996:ALS

[BWF^{+}96] Francine D. Berman, Rich Wolski, Silvia Figueira, Jennifer Schopf, and Gary Shao. Application-level scheduling on distributed heterogeneous networks. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
BERMAN/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Bode:1996:PCH

[BXC96] Paul W. Bode, Guohong Xu, and Renyue Cen. A parallel cosmological hydrodynamics code. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL http:

//www.supercomp.org/sc96/
proceedings/SC96PROC/BODE/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Chang:1996:LLC

[CCHv96] Chi-Chao Chang, Grzegorz Czajkowski, Chris Hawblitzel, and Thorsten von Eicken. Lowlatency communication on the IBM RISC System/6000 SP. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        CHANG/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Casanova:1996:NNS

[CD96] Henri Casanova and Jack Dongarra. NetSolve: A network server for solving computational science problems. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        CASANOVA/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Carrillo:1996:DLS

[CHPW96] Alex R. Carrillo, David A. Horner, John F. Peters, and John E. West. Design of a large scale discrete element soil model for high performance computing systems. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/

CARRILLO/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Chen:1996:PHM

[CSA96] Cheng Che Chen, Jaswinder Pal Singh, and Russ B. Altman. Parallel hierarchical molecular structure estimation. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
CHEN/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Clement:1996:NPM

[CSC96] Mark J. Clement, Michael R.

Steed, and Phyllis E. Crandall. Network performance modeling for PVM clusters. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
CLEMENT/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Chen:1996:PMP

[CWwK^{+}96] Ying Chen, Marianne Winslett, Szu wen Kuo, Yong Cho, Mahesh Subramaniam, and Kent Seamons. Performance modeling for the Panda array I/O library. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL http:

//www.supercomp.org/sc96/
proceedings/SC96PROC/YING/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Dincer:1996:BWW

[DF96a] Kivanc Dincer and Geoffrey C. Fox. Building a world-wide virtual machine based on Web and HPCC technologies. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        DINCER/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Ding:1996:CDA

[DF96b] Hong Q. Ding and Robert D. Ferraro. Climate data assimilation on a massively parallel supercomputer. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL http:

                                        //www.supercomp.org/sc96/
                                        proceedings/SC96PROC/DING/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

delCorral:1996:IEB

[dL96] Anna M. del Corral and Jose M. Llaberia. Increasing the effective bandwidth of complex memory systems in multivector processors. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        CORRAL/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Evangelinos:1996:CPM

[EK96] Constantinos Evangelinos and George Em Karniadakis. Communication patterns and models in prism: A spectral element-Fourier parallel Navier–Stokes solver. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
KARNIA/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Foster:1996:MCH

[FGKT96] Ian Foster, Jonathan Geisler, Carl Kesselman, and Steven Tuecke. Multimethod communication for high-performance metacomputing applications. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
FOSTER1/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Foster:1996:DSB

[FKKC96] Ian Foster, David R. Kohr, Jr., Rakesh Krishnaiyer, and Alok Choudhary. Double standards: Bringing task parallelism to HPF via the message passing interface. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        FOSTER2/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Fu:1996:SLF

[FY96] Cong Fu and Tao Yang. Sparse LU factorization with partial pivoting on distributed memory machines. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        FU/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Garcia:1996:DDD

[GAL96] Jordi Garcia, Eduard Ayguade, and Jesus Labarta. Dynamic data distribution with control flow analysis. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        GARCIA/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Grama:1996:PHS

[GKS96] Ananth Grama, Vipin Kumar, and Ahmed Sameh. Parallel hierarchical solvers and preconditioners for boundary element methods. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/

sc96/proceedings/SC96PROC/
GRAMA/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Gallo:1996:DPS

[GMW96] Steven M. Gallo, Russ Miller, and Charles M. Weeks. The design of a portable scientific tool: A case study using SnB. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
GALLO/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Hatay:1996:PCF

[Hat96] Ferhat F. Hatay. Performance of a computational fluid dynamics code on NEC and Cray supercomputers: Beyond 10 gigaflops. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
HATAY/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Hu:1996:DPI

[HJ96] Yu Hu and S. Lennart Johnsson. A data-parallel implementation of O(N) hierarchical N-body methods. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/

                                        HU/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Hsieh:1996:DCM

[HKW96] Wilson C. Hsieh, M. Frans Kaashoek, and William E. Weihl. Dynamic computation migration in DSM systems. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        HSIEH/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Huss-Lederman:1996:ISA

[HLJJ^{+}96] Steven Huss-Lederman, Elaine M. Jacobson, Jeremy R. Johnson, Anna Tsao, and Thomas Turnbull. Implementation of Strassen’s algorithm for matrix multiplication. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        JACOBSON/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Hammond:1996:AAP

[HLT96] Steven W. Hammond, Richard D. Loft, and Philip D. Tannenbaum. Architecture and application: The performance of the NEC SX-4 on the NCAR benchmark suite. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
HAMMOND/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Haake:1996:PPL

[HSS96] Bjørn Haake, Klaus E. Schauser, and Chris Scheiman. Profiling a parallel language based on fine-grained communication. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
SCHAUSER/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Jha:1996:CPB

[JMV^{+}96] Rakesh Jha, Richard C. Metzger, Brian VanVoorst, Luiz S. Pires, Wing Au, Minesh Amin, David A. Castanon, and Vipin Kumar. The C3I parallel benchmark suite — introduction and preliminary results. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
JHA/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Kumar:1996:SPA

[KCM96] Subodh Kumar, Chun-Fa Chang, and Dinesh Manocha. Scalable parallel algorithms for interactive visualization of curved surfaces. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        MANOCHA/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Karypis:1996:PMK

[KK96] George Karypis and Vipin Kumar. Parallel multilevel k-way partitioning scheme for irregular graphs. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        KARYPIS/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Kodukula:1996:TIN

[KP96] Induprakas Kodukula and Keshav Pingali. Transformations for imperfectly nested loops. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        INDUPRAK/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Lou:1996:PAO

[LF96] John Lou and John Farrara. Performance analysis and optimization on the UCLA parallel atmospheric general circulation model code. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
LOU/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Mehringer:1996:EHP

[Meh96] Susan Mehringer. Education in high performance computing via the WWW: Designing and using technical materials effectively. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
MEHRING/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Nakamura:1996:SDC

[NIY^{+}96] Takashi Nakamura, Toshiyuki Iwamiya, Masahiro Yoshida, Yuichi Matsuo, and Masahiro Fukuda. Simulation of the 3 dimensional cascade flow with numerical wind tunnel (NWT). In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
NAKAMURA/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Ogawa:1996:OOM

[OM96] Hirotaka Ogawa and Satoshi Matsuoka. OMPI: Optimizing MPI programs using partial evaluation. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        OGAWA/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Park:1996:SDA

[PC96] Chan-Ik Park and Tae-Young Choe. Striping in disk array RM2 enabling the tolerance of double disk failures. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        CIPARK/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Plimpton:1996:TDS

[PHA^{+}96] Steve Plimpton, Bruce Hendrickson, Steve Attaway, Jeff Swegle, Courtenay Vaughan, and Dave Gardner. Transient dynamics simulations: Parallel algorithms for contact detection and smoothed particle hydrodynamics. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        PLIMPTON/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Park:1996:VMV

[PSS96] Yoonho Park, Ridgway Scott, and Stuart Sechrest. Virtual memory versus file interfaces for large, memory-intensive scientific applications. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
YOONHO/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Petrini:1996:MAR

[PV96] Fabrizio Petrini and Marco Vanneschi. Minimal adaptive routing with limited injection on toroidal k-ary n-cubes. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
PETRINI/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Snell:1996:AMH

[SG96] Quinn O. Snell and John L.

Gustafson. An analytical model of the HINT performance metric. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
SNELL/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Subhlok:1996:IJM

[SGS96] Jaspal Subhlok, Thomas Gross, and Takashi Suzuoka. Impact of job mix on optimizations for space sharing schedulers. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        SUBHLOK/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Sarin:1996:PPE

[SS96] Vivek Sarin and Ahmed Sameh. Parallel preconditioners for elliptic PDEs. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        SARIN/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Teng:1996:SHS

[TWL96] Ansel Teng, Scott Whitman, and Meemong Lee. STREN: A highly scalable parallel stereo terrain renderer for planetary mission simulations. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        TENG/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Viswanathan:1996:CDS

[VL96] Guhan Viswanathan and James R. Larus. Compiler-directed shared-memory communication for iterative parallel applications. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
VISWA/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Waheed:1996:MET

[WRH96] Abdul Waheed, Diane T. Rover, and Jeffrey K. Hollingsworth. Modeling, evaluation, and testing of Paradyn Instrumentation System. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
WAHEED/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Wang:1996:RPP

[WSC96] Lei Wang, James M. Stichnoth, and Siddhartha Chatterjee. Runtime performance of parallel array assignment: An empirical study. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
CHATTER/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Yuan:1996:CCA

[YMG96] Xin Yuan, R. Melhem, and R. Gupta. Compiled communication for all-optical TDM networks. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL http:

                                        //www.supercomp.org/sc96/
                                        proceedings/SC96PROC/YUAN/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Zagha:1996:PAU

[ZLTI96] Marco Zagha, Brond Larson, Steve Turner, and Marty Itzkowitz. Performance analysis using the MIPS R10000 performance counters. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        ZAGHA/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Zaki:1996:PDM

[ZOPL96] M. J. Zaki, M. Ogihara, S. Parthasarathy, and W. Li. Parallel data mining for association rules on shared-memory multi-processors. In ACM [ACM96], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        ZAKI/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.